3. X-ray diffraction studies show that the edge length of the unit cell in crystal- line sodium chloride is 5.638 A. Calculate the internuclear Na*-Cl distance. Calculate the density of crystalline sodium chloride. Compare your answer with the experimental value obtained from the Handbook of Chemistry and Physics.
Answer #1

Hi,

NaCl crystallizes as an interpenetrated fcc packing of Na+ and Cl- ions.

Hence, edge length = 2*rcation + 2*ranion

=> Since interionic Na+--Cl- distance = rcation + ranion? = edge length / 2 = 5.638/2 = 2.819 \AA

Density of NaCl lattice will be: mass of all atoms in one unit cell / volume of unit cell

Since 1 unit cell of NaCl consists 4 Na+ and 4 Cl- ions,

mass of one unit cell = 4*23 + 4*35.5 = 234 atomic unit = 234*1.66*10-24 grams

Volume of 1 unit cell (since its a cubic lattice) = (edge length)3

= (5.638*10-8 cm)3 = 179.22*10-24 cm3  

Hence the density will be = 234*1.66/179.22 g/cm3  

= 2.167 g/cm3

reported value from literature = 2.165 g/cm3.

The value calculated here is <0.1 % away from the actual value !